Circular No. 8203

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003gx and 2003hy T. Matheson, P. Challis, and R. Kirshner, Center for Astrophysics,

report that a spectrum (range 370–750 nm) of SN 2003gx (cf. IAUC 8178), obtained by M. Calkins on Sept. 18.42 UT with the Mt. Hopkins 1.5-m telescope (+ FAST), shows it to be a type-II supernova. Adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 10203 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is 6300 km/s. A spectrum of SN 2003hy (cf. IAUC 8200), obtained by Calkins on Sept. 18.28, shows an object with a blue continuum and narrow (FWHM ≈ 300 km/s) P-Cyg lines of hydrogen. Adopting the NED recession velocity of 7355 km/s for the host galaxy, the expansion velocity derived from the minimum of the Hβ line is only 1000 km/s. The broader component of the Balmer lines often seen in type-IIn supernovae is not readily apparent. The spectrum of SN 2003hy is similar to the spectra of the type-IIn supernovae 1997bs (Van Dyk et al. 2000, PASP 112, 1532) and 2001ac (IAUC 7597). SN 2003hy may be a super-outburst of a luminous blue variable (LBV) rather than a true supernova, but the reported magnitude at discovery is brighter than that of a typical LBV outburst.

SUPERNOVAE 2003ib, 2003ic, 2003id N. Morrell and M. Hamuy, Carnegie Observatories, report that a spec-

trum (range 400–730 nm) of SN 2003ib (cf. IAUC 8201), obtained on Sept. 18.17 UT with the Las Campanas Dupont 2.5-m telescope (+ Modular Spectrograph), shows it to be a young type-II supernova characterized by a blue continuum and Hα, Hβ, and He I 5876 lines exhibiting P-Cyg profiles. A redshift of ≈ 0.024 ± 0.001 is derived, based on the maximum of the Hα emission; assuming this redshift, the minimum of the Hβ line yields an expansion velocity of −7700 km/s. A spectrum of SN 2003ic (cf. IAUC 8201), obtained on Sept. 18.25, shows it to be a type-Ia supernova; the spectrum is similar to that of SN 1994D taken 5 days before maximum. The minimum of the Si II 635.5-nm line yields an expansion velocity of −12200 km/s (adopting the NED host-galaxy recession velocity of 16690 km/s). A spectrum of SN 2003id (cf. IAUC 8201), obtained on Sept. 18.33, shows it to be a type-Ic supernova. The spectrum, which has strong absorptions centered at 491 nm (possibly Fe II blends) and 607 nm together with a P-Cyg profile whose minimum is centered at 645 nm, bears some resemblance to that of SN 1994I taken 4 days before maximum (Filippenko 1997, ARAA 35, 309).

2003 September 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green